I need like a lesson on inverse operations for a questions like -2+8n=40
Answer by zeynep(43)

Addition and subtraction are inverse operations.
Multiplication and division are inverse operations.
-2 + 8n = 40
Since addition and subtration are inverse operations, let's add +2 to both sides.
+2-2+8n = 40
+2-2 makes 0.
0 +8n = 40
8n = 40
n is multiplied by 8. To leave n alone we should DIVIDE both sides by 8 since multiplication and division are inverse operations.
8n/8 = 40/8
n = 5